OpenStudy (aryana_maria2323):

For f(x) = esin(x) use your graphing calculator to find the number of zeros for f '(x) on the closed interval [0, 2π]. a.1 b. 2 c. 3 d. 4

OpenStudy (aryana_maria2323):

Can you please help? @123AB456C @tkhunny @zepdrix

OpenStudy (tkhunny):

It's not clear what is wanted. Your level of study seems to require you to learn a little LaTex in order to communicate a little better. Do you mean \(f(x) = e^{\sin(x)}\) You have to find a way to express what you need.

OpenStudy (aryana_maria2323):

Yes that is exactly what I mean.

OpenStudy (aryana_maria2323):

Can you help? @zepdrix

OpenStudy (anonymous):

it says use a calculator right?

OpenStudy (anonymous):

\[f(x)=e^{\sin(x)}\] \[f'(x)=\cos(x)e^{\sin(x)}\] by the chain rule

OpenStudy (anonymous):

\[e^{\sin(x)}\] is never zero, so \[\cos(x)e^{\sin(x)}\] is zero only where cosine is zero

OpenStudy (aryana_maria2323):

Yes. I put it in a graph but I don't understand this part [0,2pie]

OpenStudy (aryana_maria2323):

Why are you putting cos in front of it?

OpenStudy (anonymous):

for which two values of \(x\) in \([0,2\pi]\) is cosine equal to zero?

OpenStudy (anonymous):

it does say " find the number of zeros for f '(x)" right? so step one is to find \(f'\)

OpenStudy (aryana_maria2323):

Yes, Okay I will find the that. One second.

OpenStudy (anonymous):

...

OpenStudy (aryana_maria2323):

Okay so when I put that into the graph I got 1

OpenStudy (anonymous):

whoa hold the phone what did you get for the derivative

OpenStudy (anonymous):

just asking because 1 is certainly not the answer

OpenStudy (aryana_maria2323):

I got e^sin(x) cos^2(x)-e^sin(x) sin(x)

OpenStudy (aryana_maria2323):